OpenStudy (anonymous):

You have 30 snials in your garden. After 1 week there is 45 snails in your garden. Write a differnsial equation that deskribes the snials reproduction.

OpenStudy (anonymous):

\[A(t)=30\times \frac{45}{30}^t\] where \(t\) is time in weeks after the initial observation

OpenStudy (anonymous):

probably looks nicer as \[A(t)=30\times \left(\frac{3}{2}\right)^t\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw, although your math teacher may not like it ,since i solved no differential equation probably wants something in terms of \(A(t)=30e^{kt}\) for some \(k\)